# Conghui He4, Xiaoyang Wang6, Fan Wu1

### v2

@@ 4.3 Documents
Takeaway: LLM-based agents enhance document analysis by: Complex Data Understanding (O1): Leveraging multimodal architectures like LayoutLM (xu2020layoutlm, ) and DocLLM wang2024docllm to holistically integrate textual, spatial, and visual information from complex layouts. NL-Based Interface (O2): Enabling prompt-guided interaction where natural language queries direct visual analysis, as seen in methods like VisFocus (abramovich2024visfocus, ). Semantic Operators (O3): Employing advanced RAG systems to retrieve and synthesize information from both text and images, providing context-aware answers and summaries (lewis2020retrieval, ; lei2024visdommultimodal, ). Autonomous Evolution (O4): Utilizing iterative refinement and self-correction loops in synthetic data generation to progressively improve layout and content quality (shi2025layoutcot, ; zhang2024vascar, ).
4.4 Program Language Analysis
In this section we introduce LLMs for program analysis, focusing on code-task pairs. Unlike raw source code, this data explicitly links program artifacts to their semantic properties or desired outputs, forming the basis for sophisticated analysis tasks. Examples include code paired with vulnerability labels, formal proofs, or natural language summaries. We first survey the downstream applications that this data enables and then describe the core methodologies for synthesizing this specialized data. Table IV summarizes the core applications and the synthesis techniques that enable them.
TABLE IV: Core Techniques in LLM-Based Program Analysis.
Application Synthesis Methods Key Techniques
ATP Task-Specific Generation Iterative Refinement Autoformalization Expert Iteration
Code Summarization Few-Shot Learning RAG Synthesized Data Pairs Example Retrieval
Repo-Level Completion RAG Adaptive/Iterative Retrieval Context Fusion
Automated Theorem Proving (ATP). As a frontier application, ATP requires models to generate formally verifiable mathematical proofs. Recent data synthesis pipelines combine three key techniques to reach state-of-the-art performance on challenging benchmarks. Autoformalization translates natural-language mathematical problems into formal, machine-verifiable statements in languages such as Lean4. For example, DeepSeek-Prover xin2024deepseek-prover automatically generates large-scale Lean4 proof data from mathematical competition problems. Building on this foundation, expert iteration repeatedly tasks a model to solve a wide range of problems, collects all successful solutions as new training data, and then re-trains the model to become a stronger expert in the next round. Finally, reinforcement learning with verifier feedback allows models to explore proof strategies by generating candidate proofs and receiving binary correctness signals from formal verifiers (e.g., the Lean4 compiler), enabling learning through trial and error. The combination of autoformalization, iterative self-improvement, and verifier-guided reinforcement has empowered modern systems to generate entire, correct proofs in formal languages (xin2024deepseek-prover, ; lin2025goedel-prover, ; kimina2025prover, ).
Code Summarization. By leveraging few-shot learning on synthesized code-summary pairs, LLMs significantly outperform traditional Seq2Seq models. Success relies on effective retrieval of similar code examples, using methods from token overlap to BM25, and enhancing prompts with semantic information like data flow to achieve state-of-the-art results (geng2023llmfewshot, ; ahmed2024code, ; mao2024scla, ).
Repository-Level Code Completion. LLMs are increasingly used to provide context-aware code completions by integrating information from an entire code repository. This task has progressed through a series of RAG techniques designed to resolve the so-called “Context–Latency Conundrum”, the trade-off between richer context and higher inference time. Early systems such as RepoCoder (zhang2023repocoder, ) adopt an iterative retrieval-and-generation pipeline, alternating between retrieving relevant code snippets and generating completions conditioned on the retrieved context. To reduce latency, REPOFUSE (liang2024repofuse, ) introduces efficient context fusion, combining an “analogy context” from similar code and a “rationale context” from semantic relationships, and employing Rank Truncated Generation to condense this information. More advanced frameworks such as Repoformer (wu2024repoformer, ) pursue adaptive and selective retrieval, allowing the model to self-assess when retrieval is actually beneficial and to skip it when unnecessary, thereby improving both robustness and inference speed. Pushing the conceptual boundary further, FT2Ra (guo2024ft2ra, ) proposes fine-tuning–inspired retrieval, treating each retrieval as a gradient-like update in which the retrieved document acts as a “gradient” and a tunable “learning rate” determines how strongly the retrieved information influences the evolving code representation. Together, these strategies illustrate a steady evolution toward repository-scale code completion that balances contextual richness with real-time responsiveness.
Data Synthesis for Code-Task Pairs. The ability to generate high-quality, large-scale datasets of code–task pairs is a primary driver of progress in this field. Two broad developmental paths characterize the evolution of code synthesis models. The path of intrinsic structure embeds formal grammars directly into model architectures to guarantee syntactic correctness. For example, early approaches such as SVAE (johnson2016composing, ) and SD-VAE (dai2018syntax, ) constrained the decoding process using explicit grammar rules, ensuring valid code outputs but limiting scalability due to structural rigidity. In contrast, the path of emergent capability leverages the powerful generalization ability of modern large language models through increasingly sophisticated data strategies. This trajectory advanced from scaling instruction diversity (Self-Instruct (wang2022selfinstruct, )), to systematically increasing instruction complexity (Evol-Instruct (luo2023wizardcoder, )), and finally to harvesting real-world code diversity from open-source repositories (OSS-Instruct (wei2023magicoder, )). The widespread adoption of functional-correctness metrics such as pass@@k further catalyzed the development of genuine problem-solving abilities in these models.
$\bullet$ Core Synthesis Techniques. At the heart of data synthesis lies task-specific data generation, where initial datasets are created for targeted analysis tasks. A representative example is autoformalization in automated theorem proving, which translates natural-language mathematical problems into formal, machine-verifiable statements (xin2024deepseek-prover, ; lin2025goedel-prover, ). Such task-driven pipelines provide the foundational training data needed to bootstrap powerful code models and to capture specialized reasoning patterns.
$\bullet$ Iterative Refinement and Self-Correction. Beyond one-shot generation, modern pipelines increasingly rely on iterative improvement strategies to enhance dataset quality. Expert iteration repeatedly tasks a strong model to solve large sets of problems, collects all successful solutions as new training data, and retrains the model to create progressively stronger experts (xin2024deepseek-prover, ; lin2025goedel-prover, ). Complementing this, reinforcement learning with verifier feedback allows models to explore solution spaces by generating code, receiving binary correctness signals from external tools such as compilers, and refining their outputs through trial and error. In self-debugging variants (chen2023teaching, ), models go further by using execution feedback to explain errors and autonomously correct their own outputs, dramatically improving sample efficiency and final code quality.
Takeaway: In program analysis, LLM-based agents demonstrate significant capabilities by: Complex Data Understanding (O1): Parsing entire code repositories with RAG techniques or translating natural language problems into formal code via autoformalization (wu2024repoformer, ; xin2024deepseek-prover, ). NL-Based Interface (O2): Automating the translation of requirements expressed in natural language into formal, machine-verifiable proofs or concise code summaries (xin2024deepseek-prover, ; geng2023llmfewshot, ). Semantic Operators (O3): Implementing adaptive retrieval mechanisms in repository-level code completion to find semantically relevant code snippets, going beyond simple keyword matching (zhang2023repocoder, ; liang2024repofuse, ). Autonomous Evolution (O4): Employing techniques like expert iteration, reinforcement learning with verifier feedback, and self-debugging to continuously refine their problem-solving and code generation abilities (xin2024deepseek-prover, ; chen2023teaching, ).
@@ 4.4 Program Language Analysis
TABLE V: Overview of 3D-Language Fusion and Related Techniques.
Core Technical Objective Sub-Objective Details
3D-Language Fusion Encoder-Projection Architecture Dedicated 3D encoder + projection adapter; processes raw 3D data, and projects them into LLM-consumable feature space
End-to-End Tokenization Direct tokenization of 3D features; end-to-end fusion with text; no explicit projection layer
3D-Derived Task Enhancement Geometric Reasoning Textual abstraction of geometric problems; symbolic reasoning without computation; avoids direct spatial data exposure
Textual Description Mediation Converts 3D structures into textual descriptions; uses text-only model training/inference; avoids raw 3D input
Multi-Agent Orchestration LLM orchestrates specialized models; text-based API interaction; 3D data handled internally
Cross-modal Capability Refinement Domain Adaptation Cross-modal feature alignment; domain adversarial + contrastive learning; bridges 2D–3D gap
Feature Enhancement Modules Plug-in enhancement modules; addresses 3D perception weaknesses; preserves base architecture
Figure 9: Overview of 3D Model Analysis.
@@
3D-Derived Task Enhancement Unlike 3D-Language Fusion which focuses on aligning spatial features with textual semantics, 3D-derived task enhancement emphasizes leveraging structural representations to improve downstream tasks that inherently require geometric reasoning (e.g., spatial question answering) as shown in Figure 9. The central challenge lies in $(i)$ achieving genuine structural understanding rather than relying on superficial correlations, and $(ii)$ ensuring scalability, since incorporating high-dimensional 3D signals into large models can quickly lead to prohibitive memory and computation costs.
@@
TABLE VI: Comparison of 3D Generation Methods across different paradigms.
Method Year Paradigm Uses 2D Prior
Multi-view
Consistency
Speed
Editable
Output
Key Techniques
Fantasia3D 2023 Optimization-based Medium Slow (minutes) SDS + DMTet mesh + BRDF
SweetDreamer 2023 Optimization-based ✓(tuned) High Slow (minutes) SDS + Aligned Geometric Priors
RichDreamer 2024 Optimization-based ✓(multi-prior) High Slow (minutes) SDS + Normal/Depth/Albedo Priors
Zero-1-to-3 2023 Feed-forward + Reconstruction Medium Fast (seconds) Limited View-conditioned diffusion + NeuS
Hunyuan3D-1.0 2024 Feed-forward + Reconstruction High Fast (seconds) Limited Multi-view diffusion + Sparse-view reconstructor
CraftsMan3D 2024 Native 3D Generation High Fast (seconds) 3D latent diffusion
Direct3D 2024 Native 3D Generation High Fast (seconds) D3D-VAE + D3D-DiT
MeshAnything 2024 Native 3D Generation High Fast (seconds) VQ-VAE mesh codebook + transformer
LLaMA-Mesh 2024 Native 3D Generation High Fast (seconds) LLM autoregressive mesh tokens
3D Data for LLM. Large language models have shown remarkable capabilities in reasoning over text and 2D imagery, but extending them to 3D domains introduces unique challenges. Unlike text or images, 3D data inherently encodes geometry, spatial relations, and multi-view consistency, which makes it both richer in information and more difficult to model. Recent research has therefore explored different paradigms for leveraging or directly generating 3D representations conditioned on language, aiming to unlock applications such as text-to-3D asset creation, embodied AI, and interactive design. Table VI Existing methods can be broadly grouped into three categories: optimization-based approaches that adapt powerful 2D diffusion priors for 3D generation, feed-forward pipelines that prioritize speed and determinism, and emerging native 3D generative models that attempt to learn 3D distributions directly.
@@
This line of work focuses on improving the quality and coherence of the synthesized views to enable robust reconstruction. Zero-1-to-3 liu2023zero laid a foundational stone by training a viewpoint-conditioned 2D diffusion model to generate novel views from a single image. These synthesized views, while not perfectly consistent, are sufficiently coherent to be fed into an off-the-shelf reconstruction network like NeuSwang2021neus to produce a 3D mesh, all within seconds. Hunyuan3D-1.0 yang2024hunyuan3d follows this two-stage template but scales it into a robust, practical pipeline. It employs a powerful multi-view diffusion model that can generate a set of coherent RGB views from a single image in approximately four seconds. A key innovation is its learned sparse-view reconstructor, which is specifically trained to ingest these imperfect, AI-generated views (along with the original image) and predict high-quality geometry in a single, fast forward pass. This “synthesis-then-reconstruct” design prioritizes speed and determinism for rapid asset creation.
@@
Takeaway: Reformulating 3D models using the 3D encoder and projector aligns the 3D space with the semantic space, enabling llms to understand spatial features. Complex Data Understanding (O1): Interpret point clouds and meshes using 3D encoders and projection adapters hong20233d ; xiong20253ur , which fuses 3D spatial features with natural language semantics. NL-Based Interface (O2): LLMs function as high-capacity planners huang2024protchat , decomposing tasks and orchestrating specialized models to interact with natural language. Semantic Operators (O3): LLMs perform geometric relation identification and structured reasoning via methods such as Geometry CoT wang2025large , which translates complex 3D structures into textual descriptions wang20253dsmiles . Autonomous Evolution (O4):Learn 3D distributions directly bypassing 2D priors, and use techniques like mesh codebooks chen2024meshanything to achieve inherently 3D-consistent content creation.
@@ 5 Heterogeneous Data Analysis
XMODE nooralahzadeh2024explainable assumes that heterogeneous metadata is organized within a database and leverages LLMs to decompose natural language queries into subtasks such as text-to-SQL generation and image analysis. For instance, given the query “Plot the number of paintings that depict war in each century”, XMODE generates an operation sequence comprising: $(i)$ retrieving painting metadata from the database via SQL, $(ii)$ applying VLMs to determine whether each image depicts war, and $(iii)$ producing visualization code. In practice, however, heterogeneous data is often disorganized. To address this challenge, Wang et al. wang2024interactive encode data objects using modality-specific models (e.g., BERT for text, CLIP for images) to construct an index. They further introduce a vector weight learning model that adaptively adjusts modality weights to better capture their relative importance in similarity measurement wang2024must . The user query is used to retrieve relevant data through semantic similarity for final answer generation.
@@
Takeaway: Heterogeneous data analysis agents exhibit strong capabilities in the following aspects: Complex Data Understanding (O1): Capture both structural and semantic information across diverse modalities. NL-Based Interface (O2): Facilitate cross-modal alignment, retrieval, and reasoning through natural language interaction. Semantic Operators (O3): Combine LLM-driven semantic operations with cross-modal processing capabilities. Autonomous Evolution (O4): Enable self-designing and adaptive evolution of analysis pipelines in response to changes in data sources.
@@ 6.1 Structured Data
Open World Task Adaptation. Structured data spans diverse domains with varying specifications and task requirements. LLMs trained for general-purpose analysis may fail to meet domain-specific needs. For instance, analyzing daily forms differs substantially from rigorous analyses in business or climate applications. Moreover, heterogeneity in system specifications, such as SQL dialects, poses additional adaptation challenges. To address these issues, domain adaptation through task-specific training or fine-tuning is essential. As manual annotation is often insufficient to support large-scale training, efficient automatic data generation methods warrant further exploration. In addition, techniques such as automatic query conversion and intermediate representations are crucial for bridging modality gaps, particularly between natural language and diverse SQL dialects.
@@ 6.2 Semi-Structured Data
Diverse Markup Language Formats. Markup languages encompass a wide spectrum, including HTML, XML, JSON, Markdown, LaTeX, and domain-specific standards such as Journal Article Tag Suite (JATS). However, most existing research remains concentrated on a few widely used formats, primarily including HTML, XML, and JSON which often derived from web-scraped data. This leaves domain-specific or less-studied markup languages, as well as proprietary formats in industry, underexplored. Addressing this gap requires the creation of specialized benchmarks for languages such as JATS, LaTeX, and Markdown to evaluate tasks like structure-aware question answering and format conversion. Moreover, model architectures must evolve towards structure-aware tokenization and training paradigms that explicitly leverage the hierarchical organization of markup, rather than treating markup as plain text sequences.

### v3

@@ 4.3 Documents
Takeaway (Document Analysis): (O1): Multimodal LLM architectures (xu2020layoutlm, ; wang2024docllm, ) holistically integrate textual, spatial, and visual information from complex layouts. (O2): Prompt-guided interaction where natural language queries direct visual analysis (abramovich2024visfocus, ). (O3): Employing advanced RAG systems to retrieve and synthesize information from both text and images, providing context-aware answers and summaries (lewis2020retrieval, ; lei2024visdommultimodal, ). (O4): Iterative refinement and self-correction progressively improve the quality of document analysis (shi2025layoutcot, ; zhang2024vascar, ).
Figure 9: LLM for 3D Model Analysis.
@@
TABLE III: Comparison of 3D Generation Methods across different paradigms.
Method Paradigm Uses 2D Prior
Multi-view
Consistency
Lantency
Editable
Output
Key Techniques
Fantasia3D chen2023fantasia3d Optimization-based Medium Slow (minutes)
SDS + DMTet mesh + BRDF
SweetDreamer li2023sweetdreamer Optimization-based ✓(tuned) High Slow (minutes)
SDS + Aligned Geometric Priors
RichDreamer qiu2024richdreamer Optimization-based ✓(multi-prior) High Slow (minutes)
SDS + Normal/Depth/Albedo Priors
Zero-1-to-3 liu2023zero Feed-forward + Reconstruction Medium Fast (seconds) Limited
View-conditioned diffusion + NeuS
Hunyuan3D-1.0 yang2024hunyuan3d Feed-forward + Reconstruction High Fast (seconds) Limited
Multi-view diffusion
+ Sparse-view Reconstructor
CraftsMan3D li2024craftsman3d Native 3D Generation High Fast (seconds)
3D Latent Diffusion
Direct3D wu2024direct3d Native 3D Generation High Fast (seconds)
D3D-VAE + D3D-DiT
MeshAnything chen2024meshanything Native 3D Generation High Fast (seconds)
VQ-VAE Mesh Codebook
+ Transformer
LLaMA-Mesh wang2024llama Native 3D Generation High Fast (seconds)
LLM Autoregressive Mesh Tokens
Protein-domain Task Enhancement Unlike 3D-Language Fusion which focuses on aligning spatial features with textual semantics, 3D-derived task enhancement emphasizes leveraging structural representations to improve downstream tasks that inherently require geometric reasoning (e.g., spatial question answering) as shown in Figure 9. The central challenge lies in $(i)$ achieving genuine structural understanding rather than relying on superficial correlations, and $(ii)$ ensuring scalability, since incorporating high-dimensional 3D signals into large models can quickly lead to prohibitive memory and computation costs.
@@
3D Data for LLM. Large language models have shown remarkable capabilities in reasoning over text and 2D imagery, but extending them to 3D domains introduces unique challenges. Unlike text or images, 3D data inherently encodes geometry, spatial relations, and multi-view consistency, which makes it both richer in information and more difficult to model. Recent research has therefore explored different paradigms for leveraging or directly generating 3D representations conditioned on language, aiming to unlock applications such as text-to-3D asset creation, embodied AI, and interactive design. Table III Existing methods can be broadly grouped into three categories: optimization-based approaches that adapt powerful 2D diffusion priors for 3D generation, feed-forward pipelines that prioritize speed and determinism, and emerging native 3D generative models that attempt to learn 3D distributions directly.
@@
This line of work focuses on improving the quality and coherence of the synthesized views to enable robust reconstruction. Zero-1-to-3 liu2023zero laid a foundational stone by training a viewpoint-conditioned 2D diffusion model to generate novel views from a single image. These synthesized views, while not perfectly consistent, are sufficiently coherent to be fed into an off-the-shelf reconstruction network like NeuSwang2021neus to produce a 3D mesh, all within seconds. Hunyuan3D-1.0 yang2024hunyuan3d employs a two-stage “synthesis-then-reconstruct” pipeline. It first uses a multi-view diffusion model to generate coherent RGB views from a single image in about four seconds, then applies a learned sparse-view reconstructor to predict high-quality geometry in a single forward pass—emphasizing speed and determinism for rapid asset creation.
@@
Takeaway (3D-Model Analysis): (O1): Interpret point clouds and meshes using 3D encoders and projection adapters hong20233d ; xiong20253ur to fuse 3D spatial features with natural language semantics. (O2): LLMs function as high-capacity planners huang2024protchat to decompose tasks and interact with natural language. (O3): LLMs perform geometric relation identification and structured reasoning to translate complex 3D structures into textual descriptions wang20253dsmiles . (O4): Agentic solutions for 3D model analysis needs further research.
@@ 5 Heterogeneous Data Analysis
XMODE nooralahzadeh2024explainable assumes that heterogeneous metadata is organized within a database and leverages LLMs to decompose natural language queries into subtasks such as text-to-SQL generation and image analysis. For instance, given the query “Plot the number of paintings that depict war in each century”, XMODE generates an operation sequence comprising: $(i)$ retrieving painting metadata from the database via SQL, $(ii)$ applying VLMs to determine whether each image depicts war, and $(iii)$ producing visualization code. In practice, however, heterogeneous data is often disorganized. To address this challenge, Wang et al. wang2024interactive encode data objects using modality-specific models (e.g., BERT for text, CLIP for images) to construct an index. They further introduce a vector weight learning model that adaptively adjusts modality weights to better capture their relative importance in similarity measurement wang2024must . The user query is used to retrieve relevant data through semantic similarity for final answer generation. AgenticData sun2025agenticdata is a multi-agent framework designed to facilitate cross-domain data analysis through natural language interactions. It begins by generating a logical plan in a single step, utilizing a plan validator, a plan optimizer, and a plan executor to refine and carry out the plan. The framework then applies tailored strategies depending on the data type. Additionally, it supports a variety of relational operators (e.g., filter and project) as well as semantic operators (e.g., scan and extract).
@@
Takeaway (Heterogeneous Data): (O1): Capture both structural and semantic information across diverse modalities. (O2): Facilitate cross-modal alignment, retrieval, and reasoning through natural language interaction. (O3): Combine LLM-driven semantic operations with cross-modal processing capabilities. (O4): Enable self-designing and adaptive evolution of analysis pipelines in response to changes in data sources.
6 Industry Practice
The rapid evolution of LLM/Agent technologies has driven their integration into commercial BI and data science platforms as “Copilot”-style assistants chen2021evaluating . This trend, fueled by the need to democratize data access and unlock insights from unstructured “dark data” bib:darkdata , has created a vibrant ecosystem of specialized products for diverse data modalities. Industrial applications, exemplified in Table IV, differ notably across data modalities.
TABLE IV: Representative Industry Products for LLM/Agent-driven Data Analysis.
Data Type Product Core Functionality
Structured Data Dataiku georgiev2022dataiku Unified AI platform; NL-to-Code; Automated ML pipelines.
Cognite bib:cognite_idc Industrial knowledge graph; NL query for assets & time-series.
Semi- Structured Data Tableau kvam2023chat2vis Conversational BI; NL-driven interactive visualization.
Parseu bib:parseur Agentic data extraction; Automated parsing for emails & docs.
Un- Structured Data docAnalyzer.ai bib:docanalyzer Document intelligence; Layout-aware RAG models.
viAct zhang2024towards Video analytics; Temporal reasoning for safety monitoring.
Hetero- geneous Data Dremio bib:dremio_text2sql Lakehouse query engine; Agentic cross-source query planning.
GE Predix valle2023ge Industrial IoT analytics; Cross-modal alignment of sensor data.
For structured data, the NL-based interface is dominant. Platforms like Dataiku georgiev2022dataiku embed LLMs into autonomous pipelines, where natural language orchestrates entire workflows from data prep to model deployment. In vertical domains, Cognite bib:cognite_idc leverages LLMs for semantic analysis over industrial knowledge graphs, enabling complex queries on assets and time-series data. The key “last-mile problem” remains ensuring 100% accuracy for critical queries.
In the semi-structured domain, agents tackle conversational analysis and automation. BI platforms such as Tableau kvam2023chat2vis facilitate conversational interaction for iterative visualization and insight discovery. Concurrently, extraction agents like Parseur bib:parseur automate semi-structured to structured transformations abiteboul1997querying , replacing brittle parsers with robust AI models. Their primary bottleneck is robustness against schematic or layout drifts.
For unstructured data, Retrieval-Augmented Generation lewis2020retrieval is the cornerstone. Advanced tools like docAnalyzer.ai bib:docanalyzer enhance RAG with multimodal layout-aware models to synergistically interpret text, tables, and images. In video analytics, platforms like viAct zhang2024towards apply temporal reasoning atop vision models, allowing agents to analyze event sequences for safety compliance. The industrial hurdle is scaling RAG while managing retrieval quality and permissions.
Finally, analyzing heterogeneous data remains a frontier. Lakehouse platforms like Dremio bib:dremio_text2sql employ LLMs as agentic query planners to orchestrate execution across disparate sources. Industrial IoT platforms such as GE Predix valle2023ge focus on deep cross-modal alignment, linking sensor data with textual logs for predictive analytics. The core challenge is enabling true cross-modal reasoning, a capability still in its infancy.
Placing these industry practices within the technical landscape of Figure 2 reveals the current state of maturity. While applications now span all data modalities and increasingly leverage semantic design, most commercial tools still operate within the “tool-assisted” paradigm. Development autonomy, where an agent can independently orchestrate novel analytical pipelines, remains a frontier. This gap highlights the significant hurdles in reliability and planning, setting the stage for future research.
@@ 7.1 Structured Data
Open World Task Adaptation. Structured data spans diverse domains with varying specifications and task requirements. LLMs trained for general-purpose analysis may fail to meet domain-specific needs. For instance, analyzing tables or forms differs substantially from rigorous analyses in business or climate applications. Moreover, heterogeneity in system specifications, such as SQL dialects, poses additional adaptation challenges. To address these issues, domain adaptation through task-specific training or fine-tuning is essential. As manual annotation is often insufficient to support large-scale training, efficient automatic data generation methods warrant further exploration. In addition, techniques such as automatic query conversion and intermediate representations are crucial for bridging modality gaps, particularly between natural language and diverse SQL dialects.
@@ 7.2 Semi-Structured Data
Diverse Markup Language Formats. Markup languages encompass a wide spectrum, including HTML, XML, JSON, Markdown, LaTeX, and domain-specific standards such as Journal Article Tag Suite (JATS). However, most existing research remains concentrated on a few widely used formats, primarily including HTML, XML, and JSON, which often derived from web-scraped data. This leaves domain-specific or less-studied markup languages, as well as proprietary formats in industry, underexplored. Addressing this gap requires the creation of specialized benchmarks for languages such as JATS, LaTeX, and Markdown to evaluate tasks like structure-aware question answering and format conversion. Moreover, model architectures must evolve towards structure-aware tokenization and training paradigms that explicitly leverage the hierarchical organization of markup, rather than treating markup as plain text sequences.